# Siqi Kou1∗, Jiachun Jin1, Chang Liu1∗, Ye Ma2, Jian Jia2, Quan Chen2, Peng Jiang2, Zhijie Deng1†

### v1

@@ Abstract
We introduce Orthus, an autoregressive (AR) transformer that excels in generating images given textual prompts, answering questions based on visual inputs, and even crafting lengthy image-text interleaved contents. Unlike prior arts on unified multimodal modeling, Orthus simultaneously copes with discrete text tokens and continuous image features under the AR modeling principle. The continuous treatment of visual signals minimizes the information loss for both image understanding and generation while the fully AR formulation renders the characterization of the correlation between modalities straightforward. The key mechanism enabling Orthus to leverage these advantages lies in its modality-specific heads—one regular language modeling (LM) head predicts discrete text tokens and one diffusion head generates continuous image features conditioning on the output of the backbone. We devise an efficient strategy for building Orthus—by substituting the Vector Quantization (VQ) operation in the existing unified AR model with a soft alternative, introducing a diffusion head, and tuning the added modules to reconstruct images, we can create an Orthus-base model effortlessly (e.g., within mere 72 A100 GPU hours). Orthus-base can further embrace post-training to better model interleaved images and texts. Empirically, Orthus surpasses competing baselines including Show-o and Chameleon across standard benchmarks, achieving a GenEval score of 0.58 and an MME-P score of 1265.8 using 7B parameters. Orthus also shows exceptional mixed-modality generation capabilities, reflecting the potential for handling intricate practical generation tasks.
[…]
@@ 1 Introduction
Multimodal models have shown promise in image-to-text and/or text-to-image generation, with LLaVA [34, 33], Emu2 [51], and NExT-GPT [64] as popular examples. These abilities are essential for the processing of complex real-world understanding and generation problems. Yet, existing approaches can suffer from significant modeling redundancy due to the trivial combination of specialized large models (e.g., CLIP-ViT [40], Stable Diffusion [44, 39], and LlaMa [56, 57]). Doing so also undermines the benefits brought by cross-modal learning and introduces considerable inefficiency for both training and inference.
@@
There is ongoing interest in jointly modeling interleaved image-text with a unified, compact model. One strategy is to map both images and texts to discrete tokens for simple autoregressive (AR) modeling [36, 53, 62] (see left of Figure 1). However, the image tokenizer, often equipped with a vector quantization (VQ) bottleneck, can cause inevitable information loss and easily lead to suboptimal performance on vision tasks concerning high-frequency details (e.g., OCR and human face generation). Alternatively, recent works, including Transfusion [77] and Monoformer [76] (middle of Figure 1), propose to integrate AR modeling on discrete text tokens and diffusion modeling on continuous image features within a single transformer. Nonetheless, the nature of diffusion modeling to process noisy images [21] makes the joint modeling of visual generation and understanding challenging.
@@
Figure 1: Comparison of existing unified multimodal models with Orthus. Left: Fully AR models [36, 53, 62] convert visual signals to discrete image tokens via vector quantization for joint modeling with text tokens, but this causes information loss. Middle: AR-diffusion mixed models [77, 67] perform next-token prediction for text generation and image patch denoising for image generation, but the involved noise disturbance on images makes the concurrent image-to-text and text-to-image generation challenging. Right: Orthus operates in a fully AR manner while circumventing vector quantization and noise disturbance to preserve input information and modeling flexibility.
@@
This paper proposes Orthus [^1] Orthus is a loyal two-headed guard dog in Greek mythology. to bridge the gap. Orthus conjoins lossless continuous image features and the unified, cross-modal AR modeling by decoupling diffusion from the transformer backbone. This renders the characterization of the correlation between modalities precise and straightforward. Specifically, Orthus embeds both discrete text tokens (from an off-the-shelf tokenizer) and continuous patch-wise image features (from a pre-trained variational autoencoder [25]) into the same representation space, where an AR transformer is then invoked to model the inter- and intra-modality interdependence. On top of the backbone, Orthus defines two modality-specific heads, with one as the regular language modeling (LM) head to predict discrete text tokens and the other as a novel diffusion head to craft continuous image features. During inference, Orthus autoregressively predicts the next text token or image patch according to the indication of special transition tokens.
@@
Notably, the investigation into diffusion head for superior image generation draws a striking analogy to the recent masked AR (MAR) approach [28], yet with a focus shift from image-only generation to mixed-modality one. On the other hand, our Orthus differentiates from MAR and its variant [68] in that it characterizes the correlation with fully AR formulation instead of mask-based modeling, which avoids expensive hyperparameter specification and eases the modeling of interleaved images and texts.
@@
The other important contribution of this work is a super-efficient strategy to build Orthus. Inspired by that Orthus differentiates from the representative token-based AR model Chameleon [53] only in the input embedding modules and output heads, we propose to substitute the VQ operation in Chameleon with a soft alternative and augment the model with an extra diffusion head to instantiate Orthus. We tune only the embedding modules and diffusion head (with 0.3B parameters in total) to reconstruct images on a 10k dataset to effortlessly obtain an Orthus base model. Orthus can further adopt post-training to bolster its ability to model interleaved images and text.
@@
We have performed extensive studies to evaluate Orthus. The results evidence that Orthus is substantially superior to Chameleon and Show-o [67] across multimodal understanding and generation tasks. Notably, Orthus achieves a GenEval [16] accuracy of 0.58, even surpassing specialized text-to-image models including DALL-E 2 [42] and SDXL [39]. Furthermore, on image understanding tasks like POPE [29] and GQA [22], Orthus achieves scores of 79.6 and 52.8, outperforming the performant InstructBLIP-13B [7]. We also showcase the ability of Orthus for mixed-modality generation in both zero-shot scenarios and downstream fine-tuning ones like storybook creation.
@@
  • We introduce Orthus for interleaved image-text generation. Orthus models the correlation between modalities through the AR principle and generates discrete text tokens and continuous image features with dedicated heads.
@@
  • Orthus outperforms related works like Chameleon [53] and Show-o [67] across various understanding and generation benchmarks, positioning it as a promising approach for unified multimodal modeling.
@@ 2 Related Work
Visual understanding. To enable multimodal large language models (MLLMs) to comprehend modalities beyond text, prior work has introduced methods that leverage pre-trained, modality-specific encoders [40, 27, 72, 5] to generate latent representations for each distinct modality. These representations are then projected into a pre-trained LLM’s input space through trained adapters, allowing for multimodal information alignment within the language model, understanding and reasoning are handled within the transformer backbone [34, 78, 7, 10, 4, 32, 30]. This framework allows LLMs to perform complex multimodal tasks while maintaining the language-based reasoning capabilities inherent to their architecture.
@@
Visual generation. The generation of visual content has long been a central focus within the deep learning research community [24, 17, 23, 58]. Over the past few years, research in visual generation has focused on decomposing visual signals in a more sophisticated manner and generating them iteratively. Diffusion models [45, 21, 47, 8, 44, 38, 12] transform generation into a reverse diffusion process from noise to data, gradually refining an initial noise input through a series of denoising steps. While another line of work aims to emulate the success of AR modeling from language modeling within the visual domain [37, 43, 41, 73, 49]. Specifically, images are first transformed into a sequence of vector-quantized tokens [59, 11, 55, 74], after which AR modeling is then performed on the discrete-valued token space [56]. To mitigate generation quality degradation caused by information loss during the VQ process, MAR replaces the per-token categorical distribution modeling with a diffusion procedure [28, 13]. HART [52] utilizes a hybrid tokenizer and simultaneously models the discrete per-token distribution and the continuous residual components with a diffusion model. Our proposed method generalizes MAR to cross-modality generation.
@@
Unified visual understanding and generation. To enable a model to possess both understanding and generation capabilities, one kind of approach aims to connect LLMs with multimodal adapters and diffusion decoders [50, 15, 9, 70]. However, using multiple distinct components can lead to redundancy and inefficient information use. Consequently, an increasing number of studies have sought to leverage a single transformer for unified understanding and generation. Chameleon [53] and Anole [6] use a unified token space to reason over and generate interleaved image and text sequences. LWM [35] extends its transformer’s context window to 1 million tokens to process both video and text sequences. VILA-U [66] employs contrastive learning to align the visual encoder with its textual inputs, then a unified AR next-token prediction framework is used to conduct all understanding and generation tasks. These methods all apply vector quantization to continuous visual signals to enable visual tokens, like discrete text tokens, to be trained using cross-entropy loss. As an alternative to fully AR models, some works have explored combining AR modeling with diffusion modeling. Show-o [67] unifies AR and discrete diffusion modeling for multimodal understanding and generation within one single transformer. Transfusion [77] and Monoformer [76] train one shared transformer for both discrete text autoregression and continuous image diffusion. Our proposed method circumvents the potential information loss caused by quantization and noise disturbance.
@@ 3 Preliminary
Unified multimodal modeling aims to cope with a blend of images and texts with a single compact model [36, 53, 62, 77, 67]. The model usually includes a vision autoencoder, specified with an encoder $E$ and a decoder $D$, a text tokenizer, and a transformer network [60]. The encoder $E$ is used to map the input image to a sequence of patch-wise features $V:=[{\bm{v}}_{1},\ldots,{\bm{v}}_{n}]$, ${\bm{v}}_{i}\in\mathbb{R}^{d_{v}}$ for effective information compression, where $d_{v}$ is the feature dimension and $n$ is the number of patches. The text tokenizer maps the input text into a sequence of text tokens $U:=[u_{1},\ldots,u_{m}]$ with $m$ as the sequence length. The transformer is then asked to process $U$ and $V$ simultaneously to yield meaningful outputs, which can be then detokenized as texts or decoded by $D$ to produce images. There are primarily two strategies for the learning of the transformer, detailed as follows.
Fully AR models. Observing that the AR principle excels in the generative modeling of discrete content, seminal works, including LWM [35] and Chameleon [53], propose to leverage the Vector Quantization (VQ) [59] technique to transform the continuous image features $V$ as discrete tokens to enable a fully AR modeling of the mixture of images and texts. Specifically, VQ introduces a set of $K$ codes $\{{\bm{c}}_{j}\in\mathbb{R}^{d_{v}}\}_{j=1}^{K}$ and solves the following problem for continuous-to-discrete transformation:
@@
Despite being simple, the fully AR models can suffer from information loss [31, 53], because VQ makes the transformer unable to directly look at the image features ${\bm{v}}_{i}$.
@@
AR-diffusion mixed models. Another line of unified multimodal models is AR-diffusion mixed models [77, 67, 76], which integrates diffusion modeling on images [21, 38] and AR modeling on text within a shared transformer. Take Transfusion [77] for example, its inputs are a noisy version of the image features $V$, denoted as $\bar{V}$, and the text tokens $U$. To facilitate the simultaneous processing of $\bar{V}$ and $U$, the attention mask of the transformer adopts a unique configuration—with a full-attention structure among $\bar{V}$ and a causal structure among $U$. Then, the outputs from $\bar{V}$ are directed to an output projector to predict the noise on $\bar{V}$, whereas the outcomes linked to $U$ are channeled to an LM head for next-token prediction. The training objective is the combination of AR loss and denoising loss with a balancing factor. During inference, the model operates as an AR model to generate texts and as a diffusion model to craft images, with special tokens indicating mode switching.
However, diffusion modeling inherently requires feeding noisy inputs to the model, hindering joint modeling of visual understanding (requiring clean images) and generation (requiring noisy ones). For example, Transfusion identifies a nearly 15% performance drop in image captioning when full-range noise is introduced during training [77].
@@ 4.1 Overview of Orthus
Specifically, let ${\bm{f}}_{i}$ denote the output state corresponding to the input image feature ${\bm{v}}_{i}$ and $\epsilon_{\theta}$ denote the diffusion head employed by Orthus with parameter $\theta$. The goal of $\epsilon_{\theta}$ is to predict for the next patch feature ${\bm{v}}_{i+1}$ conditioning on ${\bm{f}}_{i}$. According to common practice [21, 8], the learning objective for the diffusion head can be formalized as:
@@
where $\bm{\epsilon}\sim\mathcal{N}({\bm{0}},\mathbf{I})$ is a Gaussian noise and $t$ is a randomly sampled timestep. $\overline{\alpha}_{t}$ follows a pre-defined noise schedule [21]. In practice, $\epsilon_{\theta}$ can be a shallow multilayer perception (MLP) with three inputs (the condition ${\bm{f}}_{i}$, the scalar timestep $t$, and the noisy state). On the other hand, the LM head remains the compact linear projection followed by a softmax transformation to yield the predictive probability of the next token over the entire vocabulary.
4.2 An Efficient Strategy for Constructing Orthus
The differences between Orthus and fully AR models exist in the vision embedding module and the output head. Given that pre-training a multimodal model from scratch can be frustratingly costly but the fully AR models like LWM [35] and Chameleon [53] are readily accessible from the open-source community, we are naturally interested in deriving Orthus based on them at a minimal expense. This section elaborates on a hard-to-soft adaptation trick and an efficient training strategy to enable this.

### v2

@@ Abstract
We introduce Orthus, a unified multimodal model that excels in generating interleaved images and text from mixed-modality inputs by simultaneously handling discrete text tokens and continuous image features under the AR modeling principle. The continuous treatment of visual signals minimizes the information loss while the fully AR formulation renders the characterization of the correlation between modalities straightforward. Orthus leverages these advantages through its modality-specific heads—one regular language modeling (LM) head predicts discrete text tokens and one diffusion head generates continuous image features. We devise an efficient strategy for building Orthus—by substituting the Vector Quantization (VQ) operation in the existing unified AR model with a soft alternative, introducing a diffusion head, and tuning the added modules to reconstruct images, we can create an Orthus-base model effortlessly (e.g., within 72 A100 GPU hours). Orthus-base can further embrace post-training to craft lengthy interleaved image-text, reflecting the potential for handling intricate real-world tasks. For visual understanding and generation, Orthus achieves a GenEval score of 0.58 and an MME-P score of 1265.8 using 7B parameters, outperforming competing baselines including Show-o and Chameleon. Our code is available at https://github.com/zhijie-group/Orthus.
@@ 1 Introduction
Multimodal models have shown promise in image-to-text and/or text-to-image generation, with LLaVA Liu et al. (2024d; c), Emu2 Sun et al. (2024c), and NExT-GPT Wu et al. (2013) as popular examples. These abilities are essential for handling complex real-world understanding and generation problems. Yet, existing approaches can suffer from significant modeling redundancy due to the trivial combination of specialized large models (e.g., CLIP-ViT Radford et al. (2021), Stable Diffusion Rombach et al. (2022), and LlaMa Touvron et al. (2023a; b)). Doing so also undermines the benefits brought by cross-modal learning and introduces considerable inefficiency for both training and inference.
@@
There is ongoing interest in jointly modeling visual understanding and generation with a unified, compact model. One strategy is to map both images and texts to discrete tokens for simple autoregressive (AR) modeling Liu et al. (2024f); Team (2024); Wang et al. (2024) (left of Figure 1). However, the image tokenizer, often equipped with a vector quantization (VQ) bottleneck, can cause inevitable information loss and easily lead to suboptimal performance on vision tasks concerning high-frequency details (e.g., OCR and human face generation). Alternatively, recent works, including Transfusion Zhou et al. (2024) and Monoformer Zhao et al. (2024) (middle of Figure 1), propose to integrate AR modeling on discrete text tokens and diffusion modeling on continuous image features within a single transformer. Nonetheless, the nature of diffusion modeling to process noisy images Ho et al. (2020) makes the joint modeling of image-to-text, text-to-image, and more complicated interleaved image-text challenging.
@@
Figure 1: Comparison of existing unified multimodal models with Orthus. Left: Fully AR models Liu et al. (2024f); Team (2024); Wang et al. (2024) convert visual signals to discrete image tokens via vector quantization for joint modeling with text tokens, but this causes information loss. Middle: AR-diffusion mixed models Zhou et al. (2024); Xie et al. (2024) perform next-token prediction for text generation and image patch denoising for image generation, but the involved noise disturbance on images makes the concurrent image-to-text and text-to-image generation challenging. Right: Orthus operates in a fully AR manner while circumventing vector quantization and noise disturbance to preserve input information and modeling flexibility.
@@
This paper proposes Orthus [^1] Orthus is a loyal two-headed guard dog in Greek mythology. to bridge the gap. Orthus conjoins lossless continuous image features and the unified, cross-modal AR modeling by decoupling diffusion from the transformer backbone. This circumvents the noise disturbance and renders the characterization of the correlation between modalities straightforward, making it more suitable for interleaved image-text modeling. Specifically, Orthus embeds both discrete text tokens (from an off-the-shelf tokenizer) and continuous patch-wise image features (from a pre-trained variational autoencoder Kingma & Welling (2013)) into the same representation space, where an AR transformer is then invoked to model the inter- and intra-modality interdependence. On top of the backbone, Orthus defines two modality-specific heads, with one as the regular language modeling (LM) head to predict discrete text tokens and the other as a novel diffusion head to craft continuous image features. During inference, Orthus autoregressively predicts the next text token or image patch according to the indication of special transition tokens.
@@
Notably, the investigation into diffusion head for superior image generation draws a striking analogy to the recent masked AR (MAR) approach Li et al. (2024), yet with a focus shift from image-only generation to mixed-modality one. On the other hand, our Orthus differentiates from MAR and its variant Yang et al. (2024a) in that it characterizes the correlation with fully AR formulation instead of mask-based modeling, which avoids expensive hyperparameter specification and eases the modeling of interleaved data.
@@
The other important contribution of this work is a super-efficient strategy to build Orthus. Inspired by that Orthus differentiates from the representative token-based AR model Chameleon Team (2024) only in the input embedding modules and output heads, we propose to substitute the VQ operation with a soft alternative and augment the model with an extra diffusion head to instantiate Orthus. We tune only the embedding modules and diffusion head (with 0.3B parameters in total) to reconstruct images on a 10k dataset to effortlessly obtain an Orthus base model. Orthus-base can further adopt post-training to bolster its ability to model interleaved images and text.
@@
We have performed extensive studies to evaluate Orthus. For mixed-modality understanding and generation, Orthus outperforms the editing-specific model Instruct-pix2pix Brooks et al. (2023) and exhibits in-context learning capabilities for unseen tasks. Furthermore, Orthus demonstrates a strong ability to generate logically coherent interleaved image-text content with high relevance. For visual understanding and generation, Orthus is substantially superior to Chameleon and Show-o Xie et al. (2024) across multimodal understanding and generation tasks. Notably, Orthus achieves a GenEval Ghosh et al. (2024) accuracy of 0.58 and a POPE score of 79.6, even surpassing specialized text-to-image models SDXL Podell et al. (2023) and the performant InstructBLIP-13B Dai et al. (2023).
@@
  • We introduce Orthus for interleaved image-text generation. Orthus models the correlation between modalities through AR principle and generates discrete text tokens and continuous image features with dedicated heads.
@@
  • Compared to related works such as Chameleon Team (2024) and Show-o Xie et al. (2024), Orthus outperforms them across various visual understanding and generation benchmarks, while also demonstrating extra capabilities in mixed-modality understanding and generation, positioning it as a promising approach for unified multimodal modeling.
@@ 2 Related Work
Visual understanding. To enable multimodal large language models (MLLMs) to comprehend modalities beyond text, prior work has introduced methods that leverage pre-trained, modality-specific encoders Radford et al. (2021); Li et al. (2022); Yu et al. (2022); Chen et al. (2024) to generate latent representations for each modality. These representations are then projected into a pre-trained LLM’s input space through trained adapters, allowing for multimodal information alignment within the language model, understanding are handled within the transformer backbone Liu et al. (2024d); Zhu et al. (2023); Dai et al. (2023); Driess et al. (2023); Chen et al. (2023b); Liu et al. (2024b); Lin et al. (2024). This framework allows LLMs to perform complex multimodal tasks while maintaining the language-based reasoning capabilities inherent to their architecture.
@@
Visual generation. The generation of visual content has long been a central focus within the deep learning research community Kingma (2013); Goodfellow et al. (2014); Karras et al. (2019); Vahdat & Kautz (2020). Over the past few years, research in visual generation has focused on decomposing visual signals in a more sophisticated manner and generating them iteratively. Diffusion models Sohl-Dickstein et al. (2015); Ho et al. (2020); Song et al. (2020b); Dhariwal & Nichol (2021); Rombach et al. (2022); Peebles & Xie (2023); Esser et al. (2024) transform generation into a reverse diffusion process from noise to data, gradually refining an initial noise input through a series of denoising steps. While another line of work aims to emulate the success of AR modeling from language modeling within the visual domain Parmar et al. (2018); Razavi et al. (2019); Ramesh et al. (2021); Yu et al. (2023); Sun et al. (2024b). Specifically, images are first transformed into a sequence of vector-quantized tokens Van Den Oord et al. (2017); Esser et al. (2021); Tian et al. (2024); Yu et al. (2024), after which AR modeling is then performed on the discrete-valued token space Touvron et al. (2023a). To mitigate generation quality degradation caused by information loss during the VQ process, MAR replaces the per-token categorical distribution modeling with a diffusion procedure Li et al. (2024); Fan et al. (2024). Our proposed method generalizes MAR to cross-modality generation.
@@
Unified visual understanding and generation. To enable a model to possess both understanding and generation capabilities, one kind of approach aims to connect LLMs with multimodal adapters and diffusion decoders Sun et al. (2023); Ge et al. (2024); Ye et al. (2024). However, using multiple distinct components can lead to redundancy and inefficient information use. Consequently, recent studies seek to leverage a single transformer for unified understanding and generation. A straightforward approach is to apply vector quantization to continuous visual signals to enable visual tokens, like discrete text tokens, to be trained within a unified token space using cross-entropy loss. Representative works are LWM Liu et al. (2024e), Chameleon Team (2024), Anole Chern et al. (2024), and VILA-U Wu et al. (2024b). Alternatively, some works have explored combining AR with diffusion modeling. Show-o Xie et al. (2024) unifies AR and discrete diffusion modeling for multimodal understanding and generation within one single transformer. Transfusion Zhou et al. (2024) and Monoformer Zhao et al. (2024) train one shared transformer for both discrete text autoregression and continuous image diffusion. Our proposed method circumvents the potential information loss caused by quantization and noise disturbance.
@@ 3 Preliminary
Unified multimodal modeling aims to cope with a blend of images and texts with a single compact model Liu et al. (2024f); Team (2024); Wang et al. (2024); Zhou et al. (2024); Xie et al. (2024). The model usually includes a vision autoencoder, specified with an encoder $E$ and a decoder $D$, a text tokenizer, and a transformer network Vaswani (2017). The encoder $E$ is used to map the input image to a sequence of patch-wise features $V:=[{\bm{v}}_{1},\ldots,{\bm{v}}_{n}]$, ${\bm{v}}_{i}\in\mathbb{R}^{d_{v}}$ for effective information compression, where $d_{v}$ is the feature dimension and $n$ is the number of patches. The text tokenizer maps the input text into a sequence of text tokens $U:=[u_{1},\ldots,u_{m}]$ with $m$ as the sequence length. The transformer is then asked to process $U$ and $V$ simultaneously to yield meaningful outputs, which can be then detokenized as texts or decoded by $D$ to produce images. There are primarily two strategies for the learning of the transformer, detailed as follows.
Fully AR models. Observing that the AR principle excels in the generative modeling of discrete content, seminal works, including LWM Liu et al. (2024e) and Chameleon Team (2024), propose to leverage the Vector Quantization (VQ) Van Den Oord et al. (2017) technique to transform the continuous image features $V$ as discrete tokens to enable a fully AR modeling of the mixture of images and texts. Specifically, VQ introduces a set of $K$ codes $\{{\bm{c}}_{j}\in\mathbb{R}^{d_{v}}\}_{j=1}^{K}$ and solves the following problem for continuous-to-discrete transformation:
@@
Despite being simple, the fully AR models can suffer from information loss Liu et al. (2024a); Team (2024), because VQ makes the transformer unable to directly look at the image features ${\bm{v}}_{i}$.
@@
AR-diffusion mixed models. Another line of unified multimodal models is AR-diffusion mixed models Zhou et al. (2024); Xie et al. (2024); Zhao et al. (2024), which integrates diffusion modeling on images Ho et al. (2020); Peebles & Xie (2023) and AR modeling on text within a shared transformer. Take Transfusion Zhou et al. (2024) for example, its inputs are a noisy version of the image features $V$, denoted as $\bar{V}$, and the text tokens $U$. To facilitate the simultaneous processing of $\bar{V}$ and $U$, the attention mask of the transformer adopts a unique configuration—with a full-attention structure among $\bar{V}$ and a causal structure among $U$. Then, the outputs from $\bar{V}$ are directed to an output projector to predict the noise on $\bar{V}$, whereas the outcomes linked to $U$ are channeled to an LM head for next-token prediction. The training objective is the combination of AR loss and denoising loss with a balancing factor. During inference, the model operates as an AR model to generate texts and as a diffusion model to craft images, with special tokens indicating mode switching.
However, diffusion modeling inherently requires feeding noisy inputs to the model, hindering joint modeling of visual understanding (requiring clean images) and generation (requiring noisy ones). For example, Transfusion identifies a nearly 15% performance drop in image captioning when full-range noise is introduced during training.
@@ 4.1 Overview of Orthus
Specifically, let ${\bm{f}}_{i}$ denote the output state corresponding to the input image feature ${\bm{v}}_{i}$ and $\epsilon_{\theta}$ denote the diffusion head employed by Orthus with parameter $\theta$. The goal of $\epsilon_{\theta}$ is to predict for the next patch feature ${\bm{v}}_{i+1}$ conditioning on ${\bm{f}}_{i}$. According to common practice Ho et al. (2020); Dhariwal & Nichol (2021), the learning objective for the diffusion head can be formalized as:
@@
where $\bm{\epsilon}\sim\mathcal{N}({\bm{0}},\mathbf{I})$ is a Gaussian noise and $t$ is a randomly sampled timestep. $\overline{\alpha}_{t}$ follows a pre-defined noise schedule Ho et al. (2020). In practice, $\epsilon_{\theta}$ can be a shallow multilayer perception (MLP) with three inputs (the condition ${\bm{f}}_{i}$, the scalar timestep $t$, and the noisy state). On the other hand, the LM head remains the compact linear projection followed by a softmax transformation to yield the predictive probability of the next token over the entire vocabulary.
4.2 An Efficient Strategy for Constructing Orthus-base
The differences between Orthus and fully AR models exist in the vision embedding module and the output head. Given that pre-training a multimodal model from scratch can be frustratingly costly but the fully AR models like LWM Liu et al. (2024e) and Chameleon Team (2024) are readily accessible from the open-source community, we are naturally interested in deriving Orthus based on them at a minimal expense. This section elaborates on a hard-to-soft adaptation trick and an efficient training strategy to enable this.